OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (4, -4) with 0° ≤ θ < 360°

OpenStudy (jdoe0001):

the expression will be written like ("r value here", "angle here") pair so to find what "r" or the modulus is, keep in mind that \(\bf r^2 = x^2 + y^2\) to find the angle , well, use the trig identity of \(\bf tan(\theta) = \cfrac{y}{x}\implies tan^{-1}(tan(\theta)) = tan^{-1}\left(\cfrac{y}{x}\right)\implies \theta=tan^{-1}\left(\cfrac{y}{x}\right)\)

OpenStudy (anonymous):

so would it plugged in as =tan^-1(-4/4)?

OpenStudy (jdoe0001):

you could do that... what... do inverse trigonometric functions give anyway?

OpenStudy (anonymous):

No idea

OpenStudy (jdoe0001):

try it in your calculator... what do you get for \(\bf tan^{-1}\left(\cfrac{4}{-4}\right)\)

OpenStudy (jdoe0001):

well.... kinda got the negative in the wrong... .anyhow \(\bf tan^{-1}\left(\cfrac{-4}{4}\right)\)

OpenStudy (anonymous):

-45!

OpenStudy (anonymous):

or would it be positive?

OpenStudy (jdoe0001):

ahh... -45? ducks? gremlins? racoons? degrees!

OpenStudy (anonymous):

yes 45 degrees!

OpenStudy (jdoe0001):

so inverse trig functions give an angle, so we... -45 will be in the 4th quadrant

OpenStudy (jdoe0001):

since what you're asked is for 2 pairs.... so -45 is one.... \(\bf tan^{-1}\left(\cfrac{-4}{4}\right)=tan^{-1}(-1)\) so the other angle will be the reference angle to that, where tangent is also -1.... which means the 2nd quadrant, why? because cosine is negative and sine is positive

OpenStudy (anonymous):

so my answers on my sheet look like this. (4 , radical 2, 45°), (-4 ,radical 2, 225°) (4 ,radical 2, 135°), (-4 ,radical 2, 315°) (4 ,radical 2, 225°), (-4 ,radical 2, 45°) (4 ,radical 2, 315°), (-4 ,radical 2, 135°) how do i find the other angle and how do i know which pair of parenthesis the answers go in

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

for polars keep in mind that "r" can be negative or positive, it has no consequence however on the magnitude of "r"